OpenStudy (countrygirl1431):

Solve the following equation. 3x + 1 + 5x = 7 +15 + 7x

OpenStudy (mathstudent55):

First, combine like terms on both sides of the equal sign. Like terms are terms that have the same variable. On the left side, combine 3x and 5x. On the right side, combine 7 and 15.

OpenStudy (mathstudent55):

What do you get?

OpenStudy (mathstudent55):

\( \color{red}{3x} + 1 + \color{red}{5x} = \color{green}{7 +15} + 7x\) \(\color{red}{3x + 5x} + 1 = \color{green}{7 +15} + 7x\) What is 3x + 5x = ? What is 7 + 15 = ?

OpenStudy (triciaal):

@countrygirl1431 waiting for your input

OpenStudy (countrygirl1431):

7x and 22

OpenStudy (mathstudent55):

3 + 5 = 8, right? So 3x + 5x = 8x.

OpenStudy (mathstudent55):

\(\color{red}{8x } + 1 = \color{green}{22} + 7x\) So far we combined the like terms on the left and we combined the like terms on the right. Ok so far?

OpenStudy (countrygirl1431):

oops sorry i put thr wrong number sorry

OpenStudy (mathstudent55):

Now we want all terms with x on the left side, and all plain numbers on the right side. Since there the term 7x on the right side, we subtract 7x from both sides.